in rhombus ABCD, AB = 2x + 3 and BC = 5x find x

Answer by shahid(44)

rhombus is a four sided geometrical figure whose all sides are equal if ABCD is a rhombus and AB=2x+5 and BC=5x then AB=BC
2x+5=5x
5=5x-2x
5=3x
5/3=x
x=5/3 required value of x.